OpenStudy (anonymous):

is x+4=y and y-x=-3 parallel

OpenStudy (solomonzelman):

If you subtract x from bot sides in the first equations, you will get that the slopes are the same, and thus the lines are parallel.

OpenStudy (anonymous):

is 3x-4=y and y-3x=8

OpenStudy (solomonzelman):

subtract 3x from both sides in the first equation. Are the slopes the same ?

OpenStudy (anonymous):

parallel

OpenStudy (solomonzelman):

Yes

OpenStudy (anonymous):

is y+3=6x and -6x-y=2 parallel